OpenStudy (anonymous):

What is the product of the polynomials below? (3x + 3)(6x + 9)

OpenStudy (anonymous):

1. Solve using FOIL. 2. Add and/or subtract like terms. Do you know how to solve using foil?

OpenStudy (jessiegonzales):

Simplifying 6x + 3 = 3x + 9 Reorder the terms: 3 + 6x = 3x + 9 Reorder the terms: 3 + 6x = 9 + 3x Solving 3 + 6x = 9 + 3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3x' to each side of the equation. 3 + 6x + -3x = 9 + 3x + -3x Combine like terms: 6x + -3x = 3x 3 + 3x = 9 + 3x + -3x Combine like terms: 3x + -3x = 0 3 + 3x = 9 + 0 3 + 3x = 9 Add '-3' to each side of the equation. 3 + -3 + 3x = 9 + -3 Combine like terms: 3 + -3 = 0 0 + 3x = 9 + -3 3x = 9 + -3 Combine like terms: 9 + -3 = 6 3x = 6 Divide each side by '3'. x = 2 Simplifying x = 2

OpenStudy (anonymous):

no hahaha

OpenStudy (anonymous):

i do not know how to use foil

OpenStudy (anonymous):

(3x + 3)(6x + 9) 3x times 6x and 3x times 9, 3 times 6x and 3 times 9. That's FOIL

OpenStudy (anonymous):